OpenStudy (anonymous):

PLz help gold medal will b awarded. Part A: solve A=9/2(X+28) for X Part B: determine the value of X when A=135 Part C: solve -np-40>10 for n

OpenStudy (austinl):

Let's start with Part A: \(\displaystyle A=\frac{9}{2(x+28)}\) Is this correct?

OpenStudy (anonymous):

no

OpenStudy (austinl):

\(\displaystyle A=\frac{9}{2}(x+28)\) Better?

OpenStudy (anonymous):